The probability of an event is 1/12. Find the odds against that event.

Express your answer as a ratio using the : notation, but DO NOT include any spaces in your response.
Answer #1

Odds against the event = (1 - p)/p

here p = 1/12

odds = (11/12)/(1/12) = 11/1

odds = 11:1